# Jie Wu1∗  Haoling Li1∗  Xin Zhang2∗†‡  Jiani Guo3  Jane Luo2  Steven Liu2 Yangyu Huang2Ruihang Chu1‡Scarlett Li2Yujiu Yang1

### v1

@@ E.1 Algorithm
Table 12: Notation for SynthSmith Framework.
@@
Input: Task $q$; test inputs $\{x_{i}\}_{i=1}^{n}$; candidate solutions $\{A^{j}\}_{j=1}^{m}$.
Output: Golden solution $A_{\mathrm{golden}}$ and test suite $\mathcal{T}_{\mathrm{golden}}$, or None if verification fails.
Step 1: Consensus Voting & Weighting
for $i\leftarrow 1$ to $n$ do
 for $j\leftarrow 1$ to $m$ do
      Run $y_{i}^{j}\leftarrow A^{j}(x_{i})$
 $\hat{y}_{i}\leftarrow\arg\max_{y}\sum_{j=1}^{m}\mathbb{I}(y_{i}^{j}=y)$
 $w_{i}\leftarrow\mathrm{Weight}(x_{i})$
$\mathcal{T}_{\mathrm{candidate}}\leftarrow\{(x_{i},\hat{y}_{i},w_{i})\}_{i=1}^{n}$
Step 2: Split Candidate Set
Randomly partition $\mathcal{T}_{\mathrm{candidate}}$ into $\mathcal{T}_{\mathrm{golden}}$ and $\mathcal{T}_{\mathrm{val}}$
Step 3: Weighted Selection
for $j\leftarrow 1$ to $m$ do
 $S_{j}\leftarrow\sum_{(x_{i},\hat{y}_{i},w_{i})\in\mathcal{T}_{\mathrm{golden}}}w_{i}\cdot\mathbb{I}(A^{j}(x_{i})=\hat{y}_{i})$
$j^{\star}\leftarrow\arg\max_{j}S_{j}$
$A^{\prime}_{\mathrm{golden}}\leftarrow A^{j^{\star}}$
Step 4: Hold-out Confirmation
Compute unweighted accuracies of all $A^{j}$ on $\mathcal{T}_{\mathrm{val}}$
$j^{\dagger}\leftarrow\arg\max_{j}\text{Acc}(A^{j},\mathcal{T}_{\mathrm{val}})$
if $j^{\dagger}=j^{\star}$ then
 $A_{\mathrm{golden}}\leftarrow A^{\prime}_{\mathrm{golden}}$
 return $A_{\mathrm{golden}},\mathcal{T}_{\mathrm{golden}}$
else
 return None; // Discard task
@@
E.2 Test-Case Weighting Criteria.
We employ two distinct strategies for assigning weights to individual test cases:
@@
E.3 Error Rate for Labeling Test Outputs via Voting.
On TACO-verified, we measure a 5.27% false-positive rate under voting with 8 solutions. To assess the false-positive rate of test-output labeling, we evaluate our approach on real-world, verified datasets. Specifically, we randomly sample 500 tasks from the TACO-verified dataset, and for each task, we randomly retain 20 test cases.
@@
$n$ Labeling
(# solutions) Accuracy
4 94.39%
8 94.73%
16 95.13%
@@
E.4 Error Rate of Golden Solution
@@
Note that solution quality is strongly tied to model capability. The pass rates of the proprietary models (Qwen3-Max, Gemini2.5-pro, and GPT5-High) are presented in Table 16.
Table 16: Distribution of Proprietary LLMs’ First-Try Pass Rates on Test Cases.
Range (%) Qwen3-Max Gemini2.5-pro GPT5-High
$(0,20)$ 11.06% 9.57% 3.07%
$[20,40)$ 16.44% 14.38% 4.83%
$[40,60)$ 18.59% 17.17% 6.49%
$[60,80)$ 16.36% 15.80% 7.80%
$[80,100)$ 14.39% 14.90% 10.82%
100 23.16% 28.18% 66.98%
@@
On the TACO-verified dataset, our approach yields a 7.85% error rate in the selected golden solutions when $n=8$. The error rate further decreases as the number of rollout solutions increases. Such an error level is acceptable, indicating that the approach has the potential to be transferred to the synthetic setting.
E.5 Solvability of Generated Problem.
@@ H.1 Successful Case
Clearly, the earlier idea of "just add $\gcd(\text{last},\text{new})$" was incomplete ...
I realized each new subsequence must inherit the full score of its prefix subsequence,
and only then add the new gcd term ... (Reflection behavior)
Concretely, inheriting the full score contributes 2F(m1)2F(m-1),
and the newly formed terminal pairs contribute $\sum_{j<m}\gcd(A_{j},A_{m})\,2^{\,j-1}$ ...
Thus, I rebuild the recurrence:
$F(m)=2F(m-1)+\sum_{j=1}^{m-1}\gcd(A_{j},A_{m})\,2^{\,j-1}$ ...
this formula accounts for doubling subsequences and the fresh contributions ... (Rebuild behavior)
For brevity, denote $S(m)=\sum_{j<m}\gcd(A_{j},A_{m})\,2^{\,j-1}$ ...
@@
We maintain $U[t]=\sum 2^{\,j-1}$ for indices $j$ where $t\mid A_{j}$ ...
then $S(m)=\sum_{t\mid A_{m}}U[t]\cdot H[t]$, updated after each new element ... (Optimization behavior)
With $U[t]$ and $H[t]$, we evaluate $S(m)$ via divisors as $\sum_{t\mid A_{m}}U[t]\,H[t]$, matching the earlier definition ...
@@ H.2 Bad Case
// Instead, the solution might be to
// - for the root, the value can be any number in [1, n] that is < seg_min (if the segment is not empty)
// and > than the value of the parent (none) -> so any number in [1, seg_min-1] if the segment is not empty,
// or [1, n] if empty.
// - then for the children, they must be > than the value of the parent and < than the minimum value
// in the segment (a[i]+1, i-1) for the parent.
// But then the value for the parent is not known until we assign.
// We abandon.
@@
# vector<int> to(n, 1000000000); // to[i] = minimum value in the segment (a[i]+1, i-1)
# // We will use a segment tree for range minimum query, and we will update it as we go.
# // But we will do a DFS in increasing order of a[i] and within the same a[i] in increasing order of i.
# vector<int> order(n);
# for (int i = 0; i < n; i++) order[i] = i;
# sort(order.begin(), order.end(), [&](int i, int j) {
# if (a[i] != a[j]) return a[i] < a[j];
# return i < j;
# // segment tree for range minimum, size = 1<<19
# int size = 1;
# while (size < n) size *= 2;
# vector<int> seg(size, 1000000000);
# auto update = [&](int pos, int val) {
# pos += size;
# seg[pos] = val;
# while (pos > 1) {
# pos /= 2;
# seg[pos] = min(seg[2*pos], seg[2*pos+1]);
# auto query = [&](int l, int r) {
# int res = 1000000000;
# for (l += size, r += size; l < r; l /= 2, r /= 2) {
# if (l&1) res = min(res, seg[l++]);
# if (r&1) res = min(res, seg[--r]);
# return res;
@@
First, we find that the policy does not genuinely strive for higher rewards by producing better algorithms. Instead, it resorts to exploiting edge cases and attempting to “cheat” for partial rewards, rather than solving the problem with real solutions.
Second, we observe that some rollouts include irrelevant advertising content (e.g., Need for Speed) or adopt an editorial-style explanation of solutions similar to those on competitive programming platforms such as Codeforces. This indicates that the base model was likely exposed to raw competition-platform data during pretraining, from which such artifacts were inherited.

### v2

@@ E.1 Core Algorithm
Table 13: Notation for Our Framework.
@@
0: Task $q$; test inputs $\{x_{i}\}_{i=1}^{n}$; candidate solutions $\{A^{j}\}_{j=1}^{m}$.
0: Golden solution $A_{\mathrm{golden}}$ and test suite $\mathcal{T}_{\mathrm{golden}}$, or None.
1:Step 1: Consensus Voting & Weighting
2:for $i=1$ to $n$ do
3:  for $j=1$ to $m$ do
4:   Run $y_{i}^{j}\leftarrow A^{j}(x_{i})$
5:  end for
6:  $\hat{y}_{i}\leftarrow\arg\max_{y}\sum_{j=1}^{m}\mathbb{I}(y_{i}^{j}=y)$
7:  $w_{i}\leftarrow\mathrm{Weight}(x_{i})$
8:end for
[…]
10:Step 2: Split Candidate Set
11: Randomly partition $\mathcal{T}_{\mathrm{candidate}}$ into $\mathcal{T}_{\mathrm{golden}}$ and $\mathcal{T}_{\mathrm{val}}$
12:Step 3: Weighted Selection
13:for $j=1$ to $m$ do
14:  $S_{j}\leftarrow\sum_{(x_{i},\hat{y}_{i},w_{i})\in\mathcal{T}_{\mathrm{golden}}}w_{i}\cdot\mathbb{I}(A^{j}(x_{i})=\hat{y}_{i})$
15:end for
[…]
18:Step 4: Hold-out Confirmation
19: Compute unweighted accuracies of all $A^{j}$ on $\mathcal{T}_{\mathrm{val}}$
[…]
21:if $j^{\dagger}=j^{\star}$ then
22:  $A_{\mathrm{golden}}\leftarrow A^{\prime}_{\mathrm{golden}}$
23:  return $A_{\mathrm{golden}},\mathcal{T}_{\mathrm{golden}}$
24:else
25:  return None {Discard task}
26:end if
E.2 Test-Case Weighting Criteria
We offer two distinct strategies for assigning weights to individual test cases:
@@ E.2 Test-Case Weighting Criteria
In this work, we adopted Size-Based Weighting for data synthesis.
E.3 Validation Effectiveness Analysis
E.3.1 Error Rate for Labeling Test Outputs via Voting
On TACO-verified, we measure a 5.27% false-positive rate under voting with 8 solutions. To assess the false-positive rate of test-output labeling, we evaluate the proposed approach on real-world, verified datasets. Specifically, we randomly sample 500 tasks from the TACO-verified dataset, and for each task, we randomly retain 20 test cases.
@@ E.3.1 Error Rate for Labeling Test Outputs via Voting
$n$ (# solutions) Labeling Accuracy
44 $94.39\%$
88 $94.73\%$
$16$ $95.13\%$
@@
E.4 Solution Quality Assessment
E.4.1 Error Rate of Golden Solution
@@
Note that solution quality is strongly tied to model capability. The pass rates of the proprietary models are presented in Table 18 in the task solvability analysis section.
@@
On the TACO-verified dataset, the proposed approach yields a 7.85% error rate in the selected golden solutions when $n=8$. The error rate further decreases as the number of rollout solutions increases. Such an error level is acceptable, indicating that the approach has the potential to be transferred to the synthetic setting.
E.5 Task Solvability Analysis
E.5.1 Solvability of Generated Problem
@@ H.2 Bad Case
But then the value for the parent is not known until we assign.
We abandon.
@@
# vector<int> to(n, 1000000000); ...
@@
First, in some failure cases, we find that the policy may resort to exploiting edge cases and attempting to “cheat” for partial rewards, rather than genuinely solving the problem with robust algorithms.
Second, we observe that some rollouts include irrelevant advertising content (e.g., Need for Speed) or adopt an editorial-style explanation of solutions similar to those on competitive programming platforms. This indicates that the base model was likely exposed to raw competition-platform data during pretraining, from which such artifacts were inherited.